Status Report

Hubble Space Telescope Daily Report #2785 01/12/01

By SpaceRef Editor
January 12, 2001
Filed under ,

PERIOD COVERED:  0000Z (UTC) 01/11/01 – 0000Z (UTC) 01/12/01
Daily Status Report as of 012/0000Z
    1.1 Completed STIS/CCD 8837 (CCD Dark Monitor-Part 1)
        The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks.  The proposal completed with no reported problems.
    1.2 Completed WF/PC-2 8706 (Phoenix: "Halo/Disk” Structures In Dwarf Galaxies)
        The WF/PC-2 was used to perform hierarchical clustering scenarios for galaxy formation that predict that dwarf galaxies are the first to be formed and that they are building blocks of larger galaxies.  In this case, the spatial structure and the stellar age and metallicity distribution of the Phoenix dwarf galaxy are measured to give support or reject such scenarios.  The proposal completed with no anomalous activity.
    1.3 Completed STIS/CCD 8838 (Bias Monitor – Part 1)
        The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1×1, 1×2, 2×1, and 2×2 bin settings at gain=1, and 1×1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns.  The proposal completed nominally.
    1.4 Completed Six Sets of STIS/CCD 8562 (Probing the Large Scale Structure: Cosmic Shear Observations)
        The Space Telescope Imaging Spectrograph (CCD) was used to probe the distortion of light bundles from distant galaxies, looking at the statistical properties of the intervening inhomogeneous {dark} matter distribution.  The proposal completed nominally.
    1.5 Completed WF/PC-2/STIS/CCD 8559 (The Role of Dark Matter in Cluster Formation and Galaxy Evolution)
        The WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD) were used to observe the outer regions of massive clusters that represent transitional areas of great cosmological importance where field galaxies encounter the steep potential wells of dark matter and baryonic hot gas.  Little is known about either the dark matter profile at large radii or the morphological properties of infalling galaxies at those redshifts where strong evolution is observed in the cluster cores.  The proposal completed normally, with no reported problems.
    1.6 Completed Two Sets of WF/PC-2 8632 (A UV Atlas of Nearby Galaxies)
        The WF/PC-2 was used to perform a snapshot survey of local galaxies at UV wavelengths with the F300W filter.  The aim of the project is to build a reference UV Atlas of normal galaxies, whose optical images are well known, with the highest possible degree of information, covering all the morphological types and luminosity classes.  The proposal completed normally.
    1.7 Completed WF/PC-2 8164 (A Search for Intergalactic Globular Clusters in A1185)
        The WF/PC-2 was used to investigate the anomalously rich globular cluster populations surrounding many supergiant elliptical galaxies.  A series of deep WF/PC-2 images centered on the peak of the X-ray distribution in Abell 1185 should reveal a population of intergalactic globular cluster if it exists.  The proposal completed as planned.
    1.8 Completed STIS/MA1/MA2 8843 (Cycle 9 MAMA Dark Measurements)
        The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to perform the routine monitoring of the MAMA detector dark noise.  The proposal completed nominally.
    1.9 Completed STIS/CCD 8260 (Searching for the Hydrogen Reionization Edge of the Universe at 5HST)
        The Space Telescope Imaging Spectrograph (CCD) was used to observe one of 36 parallel orbits {4–5 fields of 5–8 orbits each} to constrain the H Lyman-edge in emission that marks the transition from a neutral to a fully ionized IGM at a predicted zion~eq5–15.  This edge is due to recombination from the H Lyman series and Lyman continuum, and can be used to constrain zion, one of the most important unknown quantities in large scale structure and cosmology.  The proposal completed nominally.
    1.10 Completed Two Sets of WF/PC-2 8816 (Cycle 9 UV Earthflats)
        The WF/PC-2 was used to obtain sequences of Earth streak flats to improve the quality of pipeline flat fields for the WFPC2 UV filter set and in order to monitor flat field stability.  There were no reported problems.
    1.11 Completed STIS/CCD/MA1 8691 (Calibrating the Cosmic Meter Stick: The Distance to the LMC Using Eclipsing Binaries)
        The Space Telescope Imaging Spectrograph (CCD and MA1) was used to measure the LMC distance which is crucial for the calibration of the Cosmic Distance Scale.  The observations completed with no reported problems.
    1.12 Completed WF/PC-2 8827 (Cycle 9 Supplemental Darks pt2/3)
        The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels.  The proposal completed nominally.
    2.1 Guide Star Acquisitions:
      Scheduled Acquisitions:  7
Successful:  7
  Scheduled Re-acquisitions:  9
Successful:  9
    2.2 FHST Updates:
Scheduled: 14
Successful: 14
    2.3 Operations Notes:
        Using ROP SR-1A, the SSR EDAC error counter was cleared four times.
        The engineering status buffer limits were updated twice per ROP DF-18A.
        As directed by an operations request, the Gyro #6 Low Mode Flight Software installation was successfully completed at 011/1447Z.  A Type 2 slew was executed nominally at 011/1452Z with Gyro #6 remaining in low mode.  A guide star acquisition was successful at 011/1509Z, acquiring in fine lock with biases comparable with those uplinked after the
installation.  Differences between the SAC-produced biases and the on-board biases were approximately 11 arc-sec/hour.  On-board biases produced during the guide star acquisition further refined the bias estimate for the new three-gyro configuration.  The Gyro #6 On-orbit Low Mode Test continues.
        A TTR was written for a period of no GCMR capability beginning at 011/1300Z.
        SSA transmitter #1 was turned on and off at 011/1620Z as directed by ROP IC-2.
        A TTR was generated when two commands had to be re-transmitted at 012/0837Z and at 012/0841Z during a 486 load uplink.
        Continuation of normal science observations and calibrations.

SpaceRef staff editor.